### v1

# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

Abstract.
Natural Language Processing is revolutionizing the way legal professionals and laypersons operate in the legal field. The considerable potential for Natural Language Processing in the legal sector, especially in developing computational tools for various legal processes, has captured the interest of researchers for years. This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses framework (Page et al., 2021), reviewing 148 studies, with a final selection of 127 after manual filtering. It explores foundational concepts related to Natural Language Processing in the legal domain, illustrating the unique aspects and challenges of processing legal texts, such as extensive document length, complex language, and limited open legal datasets. We provide an overview of Natural Language Processing tasks specific to legal text, such as Legal Document Summarization, legal Named Entity Recognition, Legal Question Answering, Legal Text Classification, and Legal Judgment Prediction. In the section on legal Language Models, we analyze both developed Language Models and approaches for adapting general Language Models to the legal domain. Additionally, we identify 15 Open Research Challenges, including bias in Artificial Intelligence applications, the need for more robust and interpretable models, and improving explainability to handle the complexities of legal language and reasoning.
copyright: acmlicensedjournal: CSURccs: General and reference Surveys and overviewsccs: Applied computing Lawccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generationccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Supervised learningccs: Computing methodologies Unsupervised learningccs: Computing methodologies Reinforcement learningccs: Computing methodologies Multi-task learningccs: Computing methodologies Machine learning approachesccs: Computing methodologies Artificial intelligenceccs: Information systems Language modelsccs: Information systems Retrieval tasks and goalsccs: Information systems Question answeringccs: Information systems Clustering and classificationccs: Information systems Summarization
1. Introduction
In recent years, advancements in Natural Language Processing (NLP) have significantly impacted the legal domain by simplifying complex tasks such as Legal Document Summarization (LDS), enhancing legal text comprehension for laypersons, and improving Legal Question Answering (LQA) and Legal Judgment Prediction (LJP) (Savelka et al., 2023; Pont et al., 2023; Janatian et al., 2023; Gesnouin et al., 2024; Chlapanis et al., 2024; Jiang et al., 2024; Li et al., 2022a). These improvements are primarily attributed to advancements in neural network architectures, such as transformer models (Vaswani et al., 2017). NLP techniques now enable machines to generate text, answer legal questions, drafting a regulation, and simulate legal reasoning, which can revolutionize legal practices (Janatian et al., 2023). Applications like contract review (Graham et al., 2023; Mamooler et al., 2022; Maroudas et al., 2022; Tuggener et al., 2020) and case prediction (Niklaus et al., 2021; Vianna et al., 2023) have been automated to a large extent, speeding up processes, reducing human error, and cutting operational costs. Additionally, the use of NLP allows lawyers and legal professionals to reduce their workload, enhance efficiency, and minimize errors in decision-making processes (Savelka et al., 2023). Despite the rapid development of NLP, challenges remain in processing lengthy documents, understanding complex language, and navigating complicated document structures (Yuen Iu and Man-Yi Wong, 2023; Vianna et al., 2023; Sovrano et al., 2021; Nguyen et al., 2024; Jiang et al., 2024; Ganguly et al., 2023; Bhattacharya et al., 2019), yet the promise of Large Language Models (LLMs) enhances the efficiency, accessibility, and precision of legal services.
Despite these advantages, the integration of NLP in the legal domain is not without challenges such as, biases and unfairness, and explainability issues. (Tamkin et al., 2021; Sheng et al., 2019; Deroy and Maity, 2023). The use of Artificial Intelligence (AI) in legal applications must follow strict standards of accuracy, fairness, and transparency, given the potential impact on clients’ lives and rights.
This survey article explores the current landscape of NLP applications within the legal domain. It discusses its potential benefits and the practical challenges it poses. NLP is a broad field covering a wide range of techniques for processing, analyzing, and understanding human language. By examining the latest advancements and applications of NLP in law, this article provides a comprehensive overview of the field. Figure 1 summarizes the scope of the survey and categorise the research into several areas: LQA, LJP, Legal Text Classification (LTC), LDS, legal Named Entity Recognition (NER), legal corpora, and legal Language Models (LMs). Each category lists relevant projects and papers and shows the work being done in each sub-field. Notably, there is comparatively less research in NER and legal corpora, whereas LDS and LQA have seen extensive research activity, with a substantial number of datasets and research contributions. This summary provides an overview of how NLP techniques are applied to various challenges in the legal domain and offers insights into future directions of AI in legal practice.
{forest}
forked edges, for tree= draw, minimum size=1cm, semithick, rounded corners, align = center, inner sep = 3mm, edge = semithick, -stealth, grow = east, l sep = 12mm, fork sep=6mm, [Legal
NLP [LQA [Huang et al. (2020); Sovrano et al. (2020); Khazaeli et al. (2021)
Zhong et al. (2020c); Askari et al. (2022)
Zhang et al. (2023b); Louis et al. (2024)
Sovrano et al. (2024); Yuan et al. (2023); Askari et al. (2024)
Chen et al. (2023); Büttner and Habernal (2024); Sovrano et al. (2021) ] ] [LJP [Luo et al. (2017); Zhong et al. (2018); Ye et al. (2018)
Yang et al. (2019); Chalkidis et al. (2019a); Medvedeva et al. (2020)
Zhong et al. (2020a); Xu et al. (2020)
Ma et al. (2021); Almuslim and Inkpen (2022); Feng et al. (2022)
Tong et al. (2024); Zhang et al. (2023a); Liu et al. (2023)
Strickson and De La Iglesia (2020); Niklaus et al. (2021)
Semo et al. (2022) ] ] [LTC [Elnaggar et al. (2018); Wei et al. (2018); Lee and Lee (2019)
Bambroo and Awasthi (2021); Song et al. (2022)
Fragkogiannis et al. (2023); Wang et al. (2022); Mamooler et al. (2022)
Chalkidis et al. (2019b); Tuggener et al. (2020); Chalkidis et al. (2021)
Papaloukas et al. (2021); Graham et al. (2023); Nguyen et al. (2021) ] ] [LDS [Gelbart and Smith (1991a, b); Moens et al. (1999); Farzindar (2004)
Polsley et al. (2016); Zhong et al. (2019); Schraagen et al. (2022)
Zhong and Litman (2022); Moro et al. (2023); Jain et al. (2024)
Liu et al. (2024); Shen et al. (2022) ] ] [Legal NER [ Dozier et al. (2010); Păis et al. (2021); Smădu et al. (2022)
Leitner et al. (2020); Au et al. (2022); Kalamkar et al. (2022) ] ] [Legal corpora [ Zheng et al. (2021); Chalkidis et al. (2022a, b)
Xiao et al. (2018); Henderson et al. (2022); Rabelo et al. (2022)
Barale et al. (2023); Niklaus et al. (2023); Park and James (2023)
Goebel et al. (2024); Östling et al. (2024); Niklaus et al. (2024) ] ] [Legal LMs [Models [ Chalkidis et al. (2020); Xiao et al. (2021)
Colombo et al. (2024); Shi et al. (2024)
Al-qurishi et al. (2022) ] ] [Methods [ Li et al. (2022a); Mamakas et al. (2022) ] ] ] ]
Figure 1. An overview of the research areas in Legal NLP and the key publications in the survey.
[…]
Hierarchical diagram of existing research in Legal NLP.
To provide a comprehensive understanding of the existing research on integrating AI within the legal domain, we present an overview of recent literature reviews, as summarized in Table 1. Most survey papers on intelligent legal systems focus either on traditional NLP technologies for specific tasks like LJP and LDS or take a broader approach but still miss certain aspects. As illustrated in Table 1, there seems to be a gap in survey papers that thoroughly examine all facets of this multidisciplinary field. Our current work aims to bridge this gap by offering a comprehensive survey of all NLP tasks, existing datasets and corpora, and LMs in the legal domain. We use ✓ to indicate papers that study the most of the existing research on each subject in legal NLP. Papers that do not address a subject receive ✗, and those that partially cover specific subjects along with their datasets or LMs are marked with –.
Table 1. Comparison of existing surveys with this work, which shows the covered topics of each survey.
References Covered subjects Published year Description
Legal Dataset
NLP Tasks in Legal
Legal LM
Large Legal Corpus
Dias et al. (2022) 2022 In this work, researchers did not only concentrate on the legal domain. Instead, they elaborated on foundational concepts of AI and NLP. They briefly explored the applications of NLP within the legal field but did not delve deeply into legal datasets, specific NLP tasks in the legal domain, or Legal LLMs. In contrast, our work provides a comprehensive analysis of all NLP tasks within the legal domain, including LQA, LDS, LTC. Additionally, our study covers large legal corpora and thoroughly examines the datasets available for each legal NLP task, areas that were not fully addressed in this work.
Sun (2023) 2023 Sun explored a limited number of research projects in the field of LLMs and the legal domain. It focused on two key NLP tasks within this field: LJP and statutory reasoning. Additionally, it examined three datasets and two LLMs relevant to these domain. The difference is that our work provides a comprehensive overview of all NLP tasks in the legal domain, as well as large legal corpora and datasets for each task, which were not fully covered in Sun’s study.
Cui et al. (2023) 2023 This paper comprehensively reviewed on LJP task. Authors analyzed 43 LJP datasets in 9 different languages. They summarized 16 evaluation metrics to evaluate three NLP tasks (text classification, text generation, and text regression) in LJP. For LMs, they explored existing Pre-trained Language Models (PLMs). Unlike this paper, our work provides comprehensive coverage of all NLP tasks in the legal domain and includes an exploration of large legal corpora.
Anh et al. (2023) 2023 This survey give a short explanation regarding challenges in legal language and how LLMs try to overcome the challenges. Then, the authors summarized six NLP tasks in the legal domain that can be addressed by LLMs. In terms of ethical and legal considerations, they discussed ‘bias and fairness’, ‘privacy and confidentiality’, ‘intellectual property’, ‘explainability and transparency’, and ‘responsible use’. The difference is that our work focuses on existing methods for all Legal NLP tasks, including LJP and LDS, available datasets for each task, and large legal corpora for pre-training and fine-tuning, whereas they concentrated on ethical and legal considerations and the impact of LLMs on NLP in legal texts.
Ganguly et al. (2023) 2023 Ganguly and his colleagues presented a comprehensive review at ECIR 2023 and discuss several key areas including the processing challenges of legal text, such as NER and sentence boundary detection. They traced the historical evolution of AI and law research from the 1980s, highlighted recent developments in NLP and IR techniques with a focus on the architectures of PLMs, and conducted a detailed survey of current issues and advancements in legal IR and NLP tasks like LDS and LJP. The review also included perspectives from the industry. In contrast to their survey, we also explore LQA and LTC tasks, which they did not cover, and examine large legal corpora, an area they did not address.
Chen et al. (2024a) 2024 This survey study focuses on LLMs in three distinct fields: finance, healthcare, and law. Although it attempts to cover all aspects of the legal domain and LLMs, the broad scope of addressing three expansive topics has resulted in a less thorough examination of many specific research cases within the legal field. In addition, it did not cover large legal corpora which are used for pre-training and fine-tuning purposes. Also, we explore methods for improving the efficiency of LLMs in legal domain.
Krasadakis et al. (2024) 2024 This survey study focuses on challenges and advances in some NLP tasks, such as NER and Relation Extraction. Unlike this survey, our study covers all legal NLP tasks, along with their corresponding datasets and large legal corpora. Additionally, we examine existing LLMs tailored for the legal domain.
The main difference between this work and previous surveys is that this survey aims to provide a more general description of all aspects of NLP tasks in the legal domain, rather than focusing solely on specific applications. The main contributions of this survey are summarized as follows: (1) This article extends beyond previous surveys by examining a broad spectrum of studies, and applications of legal NLP. By discussing datasets and large corpora in 24 languages and exploring the popular legal LMs, this survey establishes itself as an important resource in the field of legal NLP; (2) The survey offers an in-depth look at the challenges of integrating NLP with legal applications, with detailed discussions on technical solutions that tackle these issues, thereby enhancing understanding and encouraging further research in this evolving field. (3) This survey also highlights the existing research gaps in legal NLP, identifying areas that require further exploration and development, and providing a road-map for future research efforts in the legal NLP domain.
This document is organized as follows. In Section 2, Background and Foundational Concepts, we provide a detailed overview of legal language and the basic principles of NLP as they apply to the legal domain. In Section 3, we briefly explain the research methodology of this work and how we extracted the resources. In section 4, we explore various NLP tasks that are tailored for legal applications and show their unique requirements and the methodologies employed to address them. These specialized tasks leverage advanced NLP techniques to process, analyze, and extract meaningful information from legal texts, thereby facilitating more efficient and accurate legal research and decision-making. Additionally, we delve into the datasets available for training and evaluating legal NLP tasks, emphasizing their characteristics and the implications they have for model performance. Following this, in Section 5, we explore the development of LMs that has been specifically adapted for the legal field. Finally, in Section 6, we address the key challenges associated with deploying NLP technologies in legal settings, discussing both current issues and potential solutions. Since this survey contains many acronyms, Table 2 provides the list of acronyms and their meanings to make it easier to follow.
Table 2. List of Acronyms Used in the Survey.
Acronyms Meaning Acronyms Meaning
AMR Abstract Meaning Representation LJP Legal Judgment Prediction
AL Active Learning LQA Legal Question Answering
AI Artificial Intelligence LTC Legal Text Classification
BERT Bidirectional Encoder Representations from Transformers LSTM Long Short-Term Memory
Bi-GRU Bidirectional Gated Recurrent Units ML Machine Learning
Bi-LSTM Bidirectional LSTM MLM Masked Language Modeling
CanLII Canadian Legal Information Institute MGAT Multi-Graph Attention Network
CA Case Analysis ML-LJP Multi-Law aware LJP
CRF Conditional Random Field MLP Multi-Layer Perceptron
CNN Convolutional Neural Networks MPBFN Multi-Perspective Bi-Feedback Network
DL Deep Learning MTL Multi-task Learning
DAG Directed Acyclic Graph NER Named Entity Recognition
EDUs Elementary Discourse Units ORC Open Research Challenge
ECtHR European Court of Human Rights PEFT Parameter-Efficient Fine-Tuning
FSCS Federal Supreme Court of Switzerland PLM Pre-trained Language Model
GRU Gated Recurrent Unit PRISMA Preferred Reporting Items for Systematic Reviews and Meta-Analyses
GAT Graph Attention Networks QA Question Answering
GESAN Graph-Based Evidence Retrieval and Aggregation Network RNN Recurrent Neural Networks
IR Information Retrieval RL Reinforcement Learning
JEC-QA Judicial Examination of Chinese Question Answering RLHF Reinforcement Learning from Human Feedback
KG Knowledge Graph RAG Retrieval-Augmented Generation
KD Knowledge-Driven SLM Small Language Model
LLM Large Language Model SVM Support Vector Machine
LM Language Model TF-IDF Term Frequency Inverse Document Frequency
LDS Legal Document Summarization
2. Background and Foundational Concepts
In this section, we will establish an essential understanding of how NLP intersects with the legal domain. Initially, we explore the characteristics of legal documents, which are fundamental to this intersection. Legal texts are known for their complex structure, specialized words. Recognizing these attributes is essential as they significantly influence the development and application of AI technologies in legal practices.
2.1. Legal Documents
Legal documents are typically written in a descriptive language and presented in an unstructured text format. They have unique features that set them apart from other fields. These documents cover a broad array of texts essential to the functioning of the legal system. They include court filings, judgments, legislation, treaties, contracts, and legal correspondence, each serving a specific purpose and following to specific formatting and content standards that reflect legal logic and hierarchy. Legal documents are Fundamental tools for lawyers, judges, and legal scholars, facilitating case analysis, legislative review, and contract drafting. They are also essential in legal education and practice and provide the basis for legal arguments and decisions. Common examples of legal documents include case law repositories, statutory databases, and collections of legal agreements. These documents are utilized in various legal processes such as drafting legal arguments, performing legal analysis, and ensuring regulatory compliance.
2.1.1. Legal language and its characteristics
Legal language is characterized by its unique features that set it apart from everyday language, primarily due to its function in the legal system. One prominent feature is its formality, where legal language often employs a more formal vocabulary and syntax to ensure precision and avoid ambiguity (Gibbons and Turell, 2008). This formality is critical, as the precise meaning of terms can have significant legal effects. Legal texts also typically utilize passive constructions and complex sentence structures to provide detailed and comprehensive descriptions (Gibbons and Turell, 2008). These constructions help clarify responsibilities and outcomes without directly attributing actions or intentions to specific parties.
Another distinctive aspect of legal language is its reliance on specialized words and phrases. This includes terms that have specific meanings within legal contexts, archaic words that are not commonly used in everyday language, and standardized phrases that have been historically embedded in legal tradition (Gibbons and Turell, 2008). This can make legal documents less accessible to non-specialists, requiring legal professionals to interpret the content accurately.
Furthermore, legal language is heavily intertextual, meaning it frequently references other legal texts, such as statutes, regulations, and case law. This characteristic ensures that legal arguments are grounded in and supported by existing legal frameworks and previous cases. The dense use of citations and references in legal documents not only supports the arguments made but also connects the document to a broader legal discourse. Such intertextuality demands that legal professionals not only understand the texts themselves but also the broader legal context in which they operate. To illustrate the intertextuality of legal language, Figure 2 shows a sample page from the Code of Federal Regulations of the United States, extracted from the WestLaw (Westlaw, [n. d.]) website. Figure 2 displays § 40.51, Labor Certification, from 22 C.F.R. § 40.51, which is part of the Code of Federal Regulations of the United States. This section falls under Title 22, governing regulations related to foreign relations, specifically detailing the requirements and procedures for labor certification. Notably, the text includes highlighted references to other legal sources, such as INA 212(a)(5). This citation refers to the Immigration and Nationality Act, specifically section 212, which outlines various conditions for inadmissibility into the United States, under subsection (a), paragraph (5). In the “Credits” part, you can see the reference “56 FR 30422, July 2, 1991,” which points to a publication in the Federal Register. Here, “56 FR” indicates the volume number, and “30422” is the page number where the document begins. The date “July 2, 1991,” marks the publication date in the Federal Register. Additionally, a sentence in subsection (b), paragraph (1) consists of 68 words, showing the length and complexity typical of legal texts.
Figure 2. A sample page from the Code of Federal Regulations, illustrating the structured and referenced nature of legal documents.
[…]
A sample page from the Code of Federal Regulations of the United States
The specific characteristics of legal language, such as its formal vocabulary, complex syntax, and extensive use of references, present many challenges for NLP. Disambiguation titles and nested entities are other issues in legal contexts (Krasadakis et al., 2024). Disambiguation titles, such as ‘The President of USA’ requires precise identification based on contextual details like time and location. Nested entities , where titles of legislative articles refer to other laws, introduce further complexity. On the other hand, legal documents are frequently provided in non-machine-readable PDF formats, complicating data extraction and processing. These challenges highlight the need for advanced and specialized NLP solutions tailored to the legal domain.
2.2. Legal NLP
2.2.1. Introduction to Legal NLP
Legal NLP involves the application of NLP techniques to legal texts. This field is crucial as it helps automate and enhance the analysis of complex legal documents, improving efficiency and accuracy in legal research, compliance, and decision-making processes. The foundation of NLP is text, and the legal domain primarily consists of textual data (Almuslim and Inkpen, 2022), including statutes, case law, contracts, and regulations. Given the text-intensive nature of the legal field, NLP offers significant potential to transform how legal professionals interact with and utilize this vast amount of information. By leveraging advanced algorithms and Machine Learning (ML) models, Legal NLP aims to make legal texts more accessible, interpretable, and actionable.
2.2.2. Why NLP is a game-changer for the legal section
LLMs, as a part of NLP applications, such as ChatGPT (Ouyang et al., 2022), have had a significant impact since their public debut in November 2022. The legal sector, however, has been exploring the potential of AI for a longer period, applying it practically. NLP applications in the legal field are extensive, ranging from drafting client briefs and producing complex analyses from large document sets to enabling smaller firms to compete with larger ones (McGee, [n. d.]). NLP is very important in doing thorough checks when companies merge and greatly helps in legal education and learning in fast-changing fields (McGee, [n. d.]).
A notable demonstration of the capabilities of NLP applications in the legal sector occurred when GPT-4 passed the Uniform Bar Exam, underscoring the technology’s potential (Katz et al., 2024). Lawyers and law students are keenly aware of this potential, as evidenced by a LexisNexis survey released in August 2023 (LexisNexis, [n. d.]). The survey revealed that about half of all lawyers believe that LLMs will significantly transform legal practice, with nearly all (92 percent) expecting some impact. Additionally, 77 percent of respondents believe LLMs will increase the efficiency of lawyers, paralegals, and law clerks, while 63% foresee changes in how law is taught and studied. Moving forward, we will introduce specific NLP tasks in the legal domain, exploring their applications and impacts they are expected to have on the legal profession.
2.2.3. Basic foundations of NLP
The integration of NLP in the legal domain relies on several foundational techniques that enable the effective analysis and manipulation of legal texts. These fundamental methods provide the building blocks for more complex applications, transforming unstructured legal documents into structured, actionable information. By leveraging these NLP techniques, legal professionals can enhance their efficiency and accuracy, making it easier to manage and interpret vast amounts of legal data. The following section provides essential definitions of terms related to the basic foundations of NLP:
  • Tokenization: Tokenization is the process of breaking down a sequence of text into smaller units, typically words or sub-words, known as tokens. This is a fundamental step in NLP as it allows for the structured analysis of text. In the legal domain, tokenization helps in processing and understanding lengthy documents by segmenting them into manageable pieces.
  • Word Embeddings: Word Embeddings are continuous vector representations that encode the semantic meanings of words or tokens in a high-dimensional space. These representations allow models to convert individual tokens into a format suitable for neural network processing. During training, LMs develop embeddings that capture the relationships between words, such as synonyms.
  • Transformers: Transformers (Vaswani et al., 2017) are a neural network architecture designed to convert input sequences into output sequences by understanding the context and relationships among the elements of the sequence. For instance, given the input “What is the color of the sky?”, a transformer model internally processes and identifies connections among the words ‘color’, ‘sky’, and ’blue’. This understanding enables it to produce the output: “The sky is blue” (Amazon, [n. d.]). Transformer models enhance this process through a self-attention mechanism. This mechanism allows the model to analyze different parts of the sequence simultaneously instead of sequentially, helping it identify which parts are most significant.
  • PLMs: PLMs are trained on extensive corpora in a self-supervised manner, which involves tasks like recovering incomplete input sentences or auto-regressive language modeling. These models, such as Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), are initially trained on large-scale general text datasets. After pre-training, they can be fine-tuned for specific downstream tasks in the legal domain, adapting them to comprehend and process legal language for applications like document classification and information extraction.
  • Question Answering (QA): QA system is a type of NLP solution designed to answer questions posed in natural language. These systems take a user’s query and, by extracting relevant information from a dataset provide a relevant and informative response.
  • NER: NER is the task of identifying mentions of specific entities within a text that belong to predefined categories such as persons, locations, organizations, and more. NER is a fundamental component for many NLP applications, including question answering, text Summarization, and machine translation (Li et al., 2022b).
  • Information Retrieval (IR): IR involves the process of obtaining relevant information from large collections of unstructured legal texts, such as case laws, statutes, contracts, and regulations. The goal of IR is to provide users with the most relevant documents and data in response to specific queries.
  • Multi-task Learning (MTL): MTL is an approach in ML where a model is trained on multiple related tasks simultaneously or utilizes auxiliary tasks to enhance performance on a specific task. By learning from diverse tasks, MTL enables models to capture generalized and complementary knowledge, improving robustness and addressing data scarcity, particularly for low-resource tasks. MTL’s ability to share implicit knowledge across tasks often leads to performance gains and more efficient models, making it a valuable strategy for building robust and adaptable systems in NLP and other domains (Chen et al., 2024b).
  • Parameter-Efficient Fine-Tuning (PEFT): PEFT is a method for adapting PLMs that involves freezing the majority of the model’s parameters and only updating a small subset. This approach significantly reduces the computational resources and time required for fine-tuning, making it particularly effective in resource-limited scenarios, while still achieving competitive performance in tasks like text generation (Li et al., 2024a).
  • Retrieval-Augmented Generation (RAG): RAG is an advanced AI framework that enhances text creation by merging traditional information retrieval systems, with the generative power of LLMs. This integration allows the AI to access additional knowledge sources while utilizing its advanced language capabilities.
2.2.4. Key Publications and Conferences in Legal NLP
This section highlights the key journals, conferences, and workshops that serve as platforms for sharing advancements and insights in the intersection of NLP and the legal domain. These resources provide good opportunities for researchers to engage with cutting-edge research in Legal NLP.
Several leading journals focus on the intersection of AI, NLP, and the legal domain. “Artificial Intelligence and Law”, published by Springer, is a leading journal that features research articles on legal reasoning, legal IR, and legal knowledge representation. Additionally, the “Journal of Law and Information Technology”, published by Taylor & Francis , focuses on the application of information technology in law, including research AI.
Conferences significantly advance research and promote collaboration in Legal NLP. The International Conference on Artificial Intelligence and Law (ICAIL) is a biennial event that showcases advances in AI applications in the legal domain, including NLP and ML. The Conference on Legal Knowledge and Information Systems (JURIX) is an annual event that focuses on legal informatics and NLP technologies.
In Legal NLP, there are some good works which are presented in workshops. The Workshop on Automated Semantic Analysis of Information in Legal Texts delves into NLP and semantic analysis of legal texts. The Natural Legal Language Processing (NLLP) Workshop provides a platform for discussing NLP technologies tailored for legal texts and is often part of major NLP conferences. The EXplainable AI in Law (XAILA) Workshop focuses on the explainability of AI systems in legal contexts, aiming to improve transparency and trust in AI applications. The Competition on Legal Information Extraction/Entailment (COLIEE) is an annual event that challenges participants to develop innovative solutions for legal information extraction and entailment tasks.
3. Methodology
This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework (Page et al., 2021). It ensures a transparent and comprehensive assessment of research to NLP tasks within the legal sector.
3.1. Search Strategy
We performed a systematic search across two academic databases to identify relevant studies, including: Google Scholar and IEEE Xplore. Then, search queries were crafted to capture studies that focused on the application of NLP to legal tasks. The search was defined by the following queries:
  • Query 1: (“Natural Language Processing” OR “NLP”) AND (“Legal” OR “Law”)
  • Query 2: (“Legal” AND (“Named Entity Recognition” OR “NER” OR “Document Summarization” OR “Text Classification” OR “Document Classfication” OR “Judgment Prediction” OR “Question Answering” OR “Corpus” OR “Language Model”))
Our search covered publications with the following date ranges for each specific NLP task: LQA from 2020-2024, LJP from 2017-2024, LTC from 2018-2023, LDS from 2016-2024, Legal NER from 2010-2022, legal corpora from 2021-2024, and Legal LMs from 2020-2024. This ensured that we included recent advancements. Peer-reviewed journal articles and high-quality conference proceedings were prioritized, along with a secondary consideration of non-peer-reviewed sources, such as arXiv articles, where relevant.
3.2. Study Selection
A total of 148 studies were initially identified from the database search. To refine this list, we applied a manual reviewing and. This process involved:
  1. (1)
    Title and Abstract Screening: We reviewed the titles and abstracts of all retrieved studies to assess their relevance to the predefined legal NLP tasks. Studies unrelated to the core legal NLP and its tasks were excluded at this stage.
  2. (2)
    Full-Text Review: Articles that passed the initial screening went through a more detailed full-text review. This was done to confirm their relevance, quality, and alignment with the inclusion criteria. During this phase, a careful study of the literature review sections of each included research and survey was conducted. This helped ensure that the studies not only contributed original findings but also reflected a comprehensive understanding of the existing research landscape in legal NLP.
  3. (3)
    Final Selection: From the original 148 studies, 127 were included in the final list, selected based on their direct relevance to key NLP tasks within the legal sector and their methodological quality, as well as their engagement with existing literature in the field.
3.3. Eligibility Criteria
To determine which studies were included in the final synthesis, we established the following criteria:
  • Inclusion Criteria:
    • The study must focus on at least one of the following NLP tasks: Legal Question Answering, Legal Named Entity Recognition, Legal Judgment Prediction, Legal Document Summarization, Legal Text Classification, Legal Language Models, or legal corpora.
    • The study must present empirical research or significant methodological contributions to legal NLP.
    • Both peer-reviewed and non-peer-reviewed (e.g., arXiv) studies were considered if they provided valuable insights.
  • Exclusion Criteria:
    • Studies focused exclusively on unrelated areas such as information retrieval methods, pattern mining, information extraction, or similarity detection without a clear application to the specific legal NLP tasks mentioned.
    • General NLP studies without a focus on legal applications.
    • Editorials, opinion pieces, or other non-research articles.
    • Papers that did not meet basic methodological standards were not included in the final analysis.
4. NLP tasks, datasets, and large corpora in legal domains
NLP tasks in the legal domains cover a range of specialized applications that address the unique challenges and requirements of legal texts. These tasks leverage advanced NLP techniques to process, analyze, and extract meaningful information from legal documents. Legal NLP tasks help make legal research and decision-making more efficient and accurate. In this section, we explore various NLP tasks that are tailored to the legal domain and show their impact on the legal section. Additionally, we will discuss existing works and research related to each task and provide an overview of the current state of the art in this field.
Furthermore, the success of these NLP tasks heavily depends on the availability and quality of domain-specific datasets. Therefore, we will also examine the datasets commonly used in legal NLP research, exploring their characteristics and the role they play in advancing the development of NLP models for the legal domain.
4.1. Legal Question Answering
4.1.1. Task
LQA involves responding to queries about law. This task is typically done by legal professionals skilled in the relevant legal field. This process requires a comprehensive review of existing laws, careful interpretation of statutes and regulations, and the application of legal principles and past cases to the relevant facts. LQA seeks to offer precise advice on legal matters. It helps people and businesses in navigating the legal landscape and addressing legal challenges. Recently, DL has been leveraged in LQA to employ neural network models that train on extensive datasets to identify complex patterns and relationships. These models evaluate the questions posed, identify relevant legal topics, and produce suitable answers based on the patterns they have learned.
Modern ML approaches for LQA, particularly DL, rely on neural networks to understand natural language. Popular architectures include Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM), and Convolutional Neural Networks (CNN), which can be fine-tuned for specific tasks such as QA. These models generate accurate responses, adapt to new patterns, understand context, and provide relevant answers. Transformer-based models, such as BERT (Devlin et al., 2019) and ChatGPT (Ouyang et al., 2022), have proven particularly effective in NLP tasks. These models use the transformer architecture and self-attention mechanisms to learn the context of the text and understand the meaning of words. This allows them to provide relevant answers by effectively weighing the importance of different parts of the input. In the following, we will study the existing LQA works in legal domain.
Huang et al. (2020) introduce the Artificial Intelligence Law Assistant, the first Chinese LQA system that integrates a legal knowledge graph (KG) to enhance query comprehension and answer ranking. The system collects a large-scale QA corpus from an online legal forum and constructs a legal KG with over 42,000 legal concepts. It employs a knowledge-enhanced interactive attention network using Bi-LSTM and co-attention mechanisms to enrich semantic representations of QA pairs with legal domain knowledge. Additionally, it provides visual explanations for selected answers, offering users a clear understanding of the QA process.
An example of an answer retrieval system specific to Private International Law is proposed by Sovrano et al. (2020). This system integrates Term Frequency Inverse Document Frequency (TF-IDF) with deep LMs to retrieve relevant answers from an automatically generated KG of contextualized grammatical sub-trees. The KG aligns with a legal ontology based on Ontology Design Patterns, such as agent, role, event, temporal parameter, and action, to reflect the legal significance of the relationships within and between provisions.
Khazaeli et al. (2021) develop an IR-based QA system tailored to the legal domain, combining sparse vector search (BM25) and dense vector techniques (semantic embeddings) as input to a BERT-based (Devlin et al., 2019) answer re-ranking system. The system utilizes Legal GloVe and Legal Siamese BERT embeddings to enhance retrieval effectiveness. An answer finder component computes the probability of a passage answering the question using a BERT sequence binary classifier fine-tuned on question-answer pairs, improving the model’s ability to discriminate good answers.
Li et al. (2024b) introduce a retrieving-then-answering framework featuring the Graph-Based Evidence Retrieval and Aggregation Network (GESAN) to enhance QA on the Judicial Examination of Chinese Question Answering (JEC-QA) dataset (Zhong et al., 2020c). The framework leverages relevant legal knowledge by predicting question topics and retrieving legal paragraphs using BM25. GESAN aggregates the evidence and processes it along with the question and options to make accurate predictions, demonstrating improved accuracy and reasoning abilities in LQA.
Askari et al. (2022) propose a method for generating query-dependent textual profiles for lawyers to improve legal expert finding on QA platforms. Using data from the Avvo [^1] https://www.avvo.com QA forum, they focus on aspects such as sentiment, comments, and recency to create profiles. These profiles are fine-tuned using BERT models (Devlin et al., 2019), and the final aggregated score is calculated using a linear combination of profile-trained BERT models. It improves retrieval performance in the legal expert finding task.
Zhang et al. (2023b) propose a generation-based method for LQA, modeling it as a generation task to produce new, relevant answers tailored to each question. The system incorporates laws as external knowledge into the answer generation process, using a retriever to fetch applicable law articles and a generator to create answers using this knowledge. Both components are integrated into a single T5 (Raffel et al., 2020) model using MTL. It can enhance the model’s understanding and generation abilities while ensuring answers are accurate and informative.
Louis et al. (2024) present an end-to-end methodology for generating long-form answers to statutory law questions using a “retrieve-then-read” pipeline. The approach involves a retriever component that uses a bi-encoder model to fetch relevant legal provisions, followed by a generator that formulates comprehensive answers based on these provisions. The generator, an autoregressive LLM based on the Transformer architecture, employs in-context learning and PEFT to generate detailed answers. The model’s interpretability is enhanced by an extractive rationale generation strategy, ensuring responses are accompanied by verifiable justifications.
Sovrano et al. (2024) propose DiscoLQA, a discourse-based LQA system that focuses on important discourse elements like Elementary Discourse Units (EDUs) and Abstract Meaning Representations (AMRs). This approach helps the answer retriever identify the most relevant parts of the discourse, enhancing retrieval accuracy. They introduce the Q4EU dataset, containing over 70 questions and 200 answers on six European norms, demonstrating improved performance in legal QA even without domain-specific training.
Yuan et al. (2023) present a three-step approach to bridge the legal knowledge gap by creating CLIC-pages—snippets that explain technical legal concepts in layperson’s terms. They construct a legal question bank containing legal questions answered by CLIC-pages, using large-scale PLMs like GPT-3 (Brown et al., 2020) to generate machine-generated questions. The study demonstrates that machine-generated questions are more scalable and diversified, aiding in improving accessibility to legal information for non-experts.
Askari et al. (2024) propose a cross-encoder re-ranker ($CE_{FS}$) for legal answer retrieval, incorporating fine-grained structured inputs from community QA data to enhance retrieval effectiveness. They introduce the LegalQA dataset containing 9,846 questions and 33,670 lawyer-curated answers. The approach involves a two-stage ranking pipeline with a BM25 retriever followed by a re-ranker, showing that integrating question tags into the input structure can bridge the knowledge gap and improve retrieval in the legal domain.
4.1.2. Datasets
The LQA datasets are a specialized resource designed to facilitate research in the domain of legal NLP. they consist of a collection of legal questions and corresponding answers, drawn for various legal documents and case law. Most questions in the LQA datasets fall into two main categories: knowledge-driven questions (KD-questions) and case-analysis questions (CA-questions) (Zhong et al., 2020c). KD-questions are centered around the understanding of specific legal concepts, whereas CA-questions involve the analysis of actual legal cases. Both types demand advanced reasoning skills and a deep comprehension of the text, making LQA a particularly challenging task in the field of NLP.
Zhong et al. (2020c) introduce JEC-QA, a dataset with 26,365 multiple-choice questions from the National Judicial Examination of China and related websites. Each question offers four possible answers and is labeled with the type of reasoning required, such as word matching, concept understanding, numerical analysis, multi-paragraph reading, and multi-hop reasoning. This dataset poses significant challenges for QA models, highlighting the gap between machine performance and human expertise in complex legal reasoning.
Sovrano et al. (2021) present a dataset designed to evaluate automated QA systems within the domain of Private International Law. It includes 17 carefully selected questions based on key EU regulations—Rome I, Rome II, and Brussels I bis—with answers derived directly from these regulations. The questions are classified based on their specificity, allowing for nuanced analysis of context-dependency in legal reasoning. This dataset aids in assessing the performance of QA systems intended for legal professionals navigating complex cross-border legal issues.
EQUALS (Chen et al., 2023) is a large-scale annotated LQA dataset in Chinese law, containing 6,914 question-answer pairs with answers based on specific law articles. Curated by senior law students, it covers 10 collections of Chinese laws and includes annotations indicating the type of reasoning required for each question. The dataset ensures that answers are precise excerpts from relevant law articles, making it valuable for developing advanced LQA systems that can aid in legal research and decision-making.
Büttner and Habernal (2024) introduce GerLayQA, a dataset supporting LQA for laypersons in Germany, focusing on the civil-law system. It contains 21,538 real-world questions posed by laypersons, paired with expert answers from lawyers grounded in specific paragraphs of German law books. The dataset are constructed through filtering and quality assurance to ensure accuracy and relevance, making it a valuable resource for developing LQA systems that can interpret and apply German law to everyday legal inquiries.
4.2. Legal Judgment Prediction
4.2.1. Task
LJP is an important task within legal NLP, especially prevalent in civil law systems where judgments are determined based on case facts and statutory articles (Zhong et al., 2020b). This task involves predicting legal outcomes from the descriptions of cases and relevant legal statutes, and is essential in countries like France, Germany, Japan, and China (Zhong et al., 2020b). LJP has garnered significant attention from AI researchers and legal professionals due to its potential to assist judges, lawyers, and legal scholars in predicting case outcomes based on historical data (Chen et al., 2024a).
Despite its promise, LJP is a highly demanding and challenging task. It requires careful handling of natural biases in historical legal data, which can create feedback loops that amplify discrimination (Krasadakis et al., 2024). Therefore, ensuring the impartiality of predicted rulings is crucial (Krasadakis et al., 2024). Currently, LJP is primarily performed by legal experts who undergo extensive specialized training to manage the complex steps involved, such as identifying relevant law articles, defining charge ranges, and deciding penalty terms (Cui et al., 2023). Nevertheless, LJP provides substantial benefits, streamlining legal decision-making processes for both practitioners and ordinary citizens (Cui et al., 2023).
Luo et al. (2017) propose an attention-based neural network to enhance charge prediction by jointly modeling charge prediction and relevant law article extraction. They used Bidirectional Gated Recurrent Units (Bi-GRUs) to encode fact descriptions and an article extractor to identify top relevant law articles. The model employs an attention mechanism guided by context vectors to combine embeddings for prediction. Evaluations on Chinese judgment documents showed improved accuracy in predicting charges and providing relevant legal articles.
Zhong et al. (2018) introduce TopJudge, a topological MTL framework that models dependencies among subtasks in LJP, such as law article prediction, charge prediction, and penalty terms. Using a Directed Acyclic Graph (DAG), TopJudge processes subtasks in a topological order reflecting real-world legal decision-making. Evaluated on large-scale Chinese criminal case datasets, it outperformed previous models in predicting legal outcomes.
Ye et al. (2018) address the problem of Court View Generation from fact descriptions in criminal cases to enhance the interpretability of charge prediction systems and aid in automatic legal document generation. They formulated this as a text-to-text Natural Language Generation (NLG) problem, using a label-conditioned Seq2Seq model with attention to decode court views based on encoded charge labels. Their model outperformed basic Seq2Seq models in generating accurate and natural court views. This work contributes to automatic legal document generation by providing justifications for charge decisions.
Yang et al. (2019) propose a Multi-Perspective Bi-Feedback Network (MPBFN) with a Word Collocation Attention mechanism to improve LJP. MPBFN addresses the challenges of multiple subtasks and their dependencies by using a bi-feedback mechanism for forward prediction and backward verification among subtasks. The Word Collocation Attention integrates word collocation features and numerical semantics to better predict penalties. Evaluated on the CAIL-small and CAIL-big datasets (Xiao et al., 2018), their model outperformed baselines in predicting law articles, charges, and penalty terms.
Chalkidis et al. (2019a) introduce an English LJP dataset containing approximately 11.5k cases from the European Court of Human Rights (ECHR). They evaluated various neural models on this dataset, including a hierarchical version of BERT  (Devlin et al., 2019) (HIER-BERT) to handle long legal documents. Their models outperformed previous feature-based approaches in tasks like violation classification and case importance prediction. They also explored potential biases in legal predictive models using data anonymization.
Medvedeva et al. (2020) investigate using NLP tools to predict judicial decisions of the ECHR based on court proceeding texts. They employed an SVM linear classifier to predict violations of articles, achieving an average accuracy of 75%. However, when predicting future decisions based on past cases, accuracy decreased to 58–68%. The study also found that predicting outcomes based solely on judges’ surnames could achieve an average accuracy of 65%, highlighting potential biases.
Zhong et al. (2020a) introduce QAjudge, a reinforcement learning-based model designed to provide interpretable legal judgments by visualizing the prediction process. QAjudge uses a Question Net to iteratively select relevant yes-no questions about case facts, an Answer Net to provide answers, and a Predict Net to generate the final judgment. The model aims to minimize the number of questions asked. It focuses on crucial elements to ensure fairness and interpretability. Evaluated on real-world datasets, QAjudge demonstrated potential in providing reliable and transparent legal judgments.
Xu et al. (2020) propose the Law Article Distillation based Attention Network (LADAN), an end-to-end model addressing the issue of confusing charges in LJP by distinguishing similar law articles. The model uses a novel graph neural network to learn differences between confusing law articles and an attention mechanism to extract discriminative features from fact descriptions. Experiments on real-world datasets showed that LADAN improved performance over previous methods in law article prediction, charge prediction, and penalty term prediction.
Strickson and De La Iglesia (2020) present the first LJP model for UK court cases, creating a labeled dataset of UK court judgments spanning 100 years. They evaluated various ML models and feature representations, with their best model achieving an accuracy of 69%. The study demonstrated the potential of LJP for UK courts, though challenges remain due to the complexity of legal language and lack of structured public datasets.
Ma et al. (2021) introduce MSJudge, a MTL framework designed to predict legal judgments by leveraging multi-stage judicial data, including pre-trial claims and court debates. MSJudge consists of components to encode multi-stage context, model interactions among claims, facts, and debates, and predict judgments. Evaluated on a large civil trial dataset [^2] https://github.com/mly-nlp/LJP-MSJudge, MSJudge outperformed state-of-the-art baselines, enhancing trial efficiency and judgment quality.
Almuslim and Inkpen (2022) focus on LJP for Canadian appeal court cases, employing various NLP and ML methods to predict binary outcomes (‘Allow’ or ‘Dismiss’) based on case descriptions. Deep learning models using custom Word2Vec embeddings achieved the highest accuracy of 93%, significantly outperforming classical ML models. The study highlights the potential of predictive models to aid legal professionals and establishes a foundation for future research in the Canadian legal system.
Feng et al. (2022) address limitations of state-of-the-art LJP models by proposing an event-based prediction model with constraints to improve performance. The model extracts fine-grained key events from case facts and predicts judgments based on these events rather than the entire fact statement. They manually annotated a legal event dataset and introduced output constraints to guide learning. Their method effectively leverages event information and cross-task consistency constraints, enhancing LJP accuracy.
Tong et al. (2024) introduce GJudge, a graph boosting framework incorporating constraints to address shortcomings of traditional LJP methods. GJudge features a multi-perspective interactive encoder and a Multi-Graph Attention Network (MGAT) consistency expert module. The encoder merges fact descriptions with label similarity connections, while the expert module distinguishes similar labels and maintains task consistency. Testing on datasets showed that GJudge outperformed other models, including the state-of-the-art RLJP (Wu et al., 2022) model, with higher F1 scores.
Previous works mainly focus on creating accurate representations of a case’s fact description to enhance judgment prediction performance. However, these methods often overlook the practical judicial process, where human judges compare similar law articles or potential charges before making a decision. To address this gap, Zhang et al. (2023a) propose CL4LJP, a supervised contrastive learning framework to improve LJP by capturing fine-grained differences between similar law articles and charges. The framework includes contrastive learning tasks at the article, charge, and label levels, enhancing the model’s ability to model relationships between fact descriptions and labels. Experiments demonstrated that CL4LJP outperformed previous methods, proving its effectiveness and robustness.
Liu et al. (2023) propose ML-LJP, a Multi-Law aware LJP method that expands law article prediction into a multi-label classification task incorporating both charge-related and term-related articles. The approach uses label-specific representations and contrastive learning to distinguish similar definitions. A Graph Attention Network (GAT) is employed to learn interactions among multiple law articles for prison term prediction. Experiments showed that ML-LJP outperformed state-of-the-art models, particularly in prison term prediction.
4.2.2. Datasets
The LJP datasets are specialized resources designed to advance research in predicting judicial outcomes within the domain of legal NLP. These datasets are categorized into four main types: court view generation datasets, law articles, charge prediction, and prison term prediction. Court view generation datasets involve court opinions and summaries. Law Articles datasets involve the prediction of legal outcomes based on specific statutes or regulations. Charge prediction datasets are concerned with predicting the charges that should be brought against a defendant based on the case details, while prison term prediction datasets aim to estimate the likely sentence duration given the nature of the crime and the legal context. Each type of dataset presents unique challenges, demanding not only text comprehension but also the ability to apply complex legal reasoning, making LJP a particularly complex task in the field of NLP.
Court View Gen (Ye et al., 2018) is an innovative dataset containing 171,981 Chinese legal cases, each involving a single defendant and a corresponding charge, covering a total of 51 charge categories. This dataset is specifically curated to support the generation of court opinions based on charge labels. The data was sourced from publicly available legal documents within the CJO [^3] http://wenshu.court.gov.cn repository.
Niklaus et al. (2021) introduce a multilingual LJP dataset from the Federal Supreme Court of Switzerland (FSCS), containing over 85,000 cases in German, French, and Italian. The dataset is annotated with publication years, legal areas, and cantons of origin, making it suitable for NLP applications in judgment prediction.
Semo et al. (2022) introduce the first LJP dataset focused on class action cases in the United States. The dataset targets predicting outcomes based on plaintiffs’ complaints rather than court-written fact summaries, involving a rule-based extraction system to identify relevant text spans from complaints.
Almuslim and Inkpen (2022) construct a dataset for LJP within the Saskatchewan Court of Appeal. They collected and labeled 3,670 documents with case outcomes (‘allow’ or ‘dismiss’) using a two-step pattern matching and keyword-based validation, ensuring label accuracy through manual annotation. This dataset supports research in the Canadian legal system and aids in developing predictive models for legal judgments.
4.3. Legal Text Classification
4.3.1. Task
LTC is an important task within the domain of NLP that involves categorizing legal documents based on their content, a foundational aspect of building intelligent legal systems. With the exponential growth of legal documents, it has become increasingly challenging for legal professionals to efficiently locate relevant rulings in similar cases for argumentation. LTC addresses this challenge by automatically associating legal texts with predefined categories, such as criminal, civil, or administrative cases, thereby simplifying legal research and decision-making processes.
In the legal field, this process is often referred to as predictive coding, where ML algorithms are trained through supervised learning to classify documents into specific categories. The broader task of text classification in NLP involves assigning one or multiple categories to a document from a set of predefined options, and it can take various forms, including binary classification, multi-class classification, and multi-label classification. Legal document classification often falls under large multi-label text classification, where the label space can consist of thousands of potential categories, adding significant complexity to the task (Shaheen et al., 2020). This subsection explores the methodologies and advancements in this area.
DL typically requires extensive data to yield effective results, but MTL offers a potential solution to the data scarcity problem. Elnaggar et al. (2018) leverage transfer learning and MTL to perform tasks like translation and multi-label classification within legal document corpora. They utilize the MultiModel algorithm (Kaiser et al., 2017), which uses a fully convolutional sequence-to-sequence architecture integrating different modality nets. Their model processes legal texts through a unified embedding, enabling efficient task switching and promoting generalization across different legal tasks, effectively tackling data scarcity in the legal field.
Wei et al. (2018) investigate the application of CNNs for text classification in legal document review, comparing their performance with SVMs on four datasets from real legal cases. They found that CNNs perform better with larger training datasets and offer a more stable growth trend compared to SVMs. The study highlighted challenges such as sequence length limitations and the need for explainability in predictive analysis, suggesting improvements like integrating annotated sentences with full text to enhance sentence relevance identification.
Lee and Lee (2019) focuse on legal document classification in the Korean language and compare three different DL approaches: CNN with ASCII encoding, CNN with Word2Vec embeddings, and RNN with Word2Vec embeddings. The classification models are used to classify case data into civil, criminal, and administrative. Using a dataset of nearly 60,000 past case documents, the study finds that the RNN model with Word2Vec embedding achieves the highest classification accuracy.
Bambroo and Awasthi (2021) introduce an architecture that integrates long attention mechanisms with a distilled BERT model pre-trained on legal domain-specific corpora. Their model employs a combination of local windowed attention and task-motivated global attention to handle inputs up to eight times longer than standard BERT models. The architecture, based on DistilBERT (Sanh et al., 2020) and incorporating LongformerSelf-Attention, is optimized for legal document classification, outperforming fine-tuned BERT and other transformer-based models in both speed and effectiveness.
Song et al. (2022) present a deep learning-based system built on top of RoBERTa (Liu et al., 2019) for multi-label legal document classification. They enhance the model with domain-specific pre-training, a label-attention mechanism, and MTL to improve classification accuracy, particularly for low-frequency classes. The label-attention mechanism uses label embeddings to bridge the semantic gap between samples and class labels, addressing class imbalance issues.
Fragkogiannis et al. (2023) propose a method to improve classification of pages within lengthy documents by leveraging sequential context from preceding pages. They enhance the input to pre-trained models like BERT (Devlin et al., 2019) by appending special tokens representing the predicted page type of the previous page, enabling more context-aware classification without modifying the model architecture. Experiments on legal datasets demonstrate improvements compared to non-recurrent setups.
Wang et al. (2022) introduce a Document-to-Graph Classifier to classify legal documents based on facts and reasons rather than topics. They extract key entities and represented legal documents using four distinct relation graphs capturing different aspects of entity relationships. A GATs (Veličković et al., 2018) is used to learn document representations from the combined graph, improving classification by focusing on factual content.
Mamooler et al. (2022) propose an active learning pipeline for fine-tuning PLMs for LTC. They address challenges of specialized vocabulary and high annotation costs. Their method involves continued pre-training of RoBERTa (Liu et al., 2019) on legal texts, knowledge distillation using a pre-trained sentence transformer, and an efficient initial sampling strategy by clustering unlabeled data. This approach reduces the number of labeling actions required and improves efficiency in adapting models to LTC tasks.
4.3.2. Datasets
LTC datasets are characterized by their domain-specific vocabulary and multi-label nature, requiring models to interpret complex legal texts and categorize them into single or multiple legal themes.
Chalkidis et al. (2019b) release EURLEX57K, a dataset containing 57,000 EU legislative documents from the EUR-LEX portal444lex.europa.eu/, annotated with EUROVOC [^5] http://eurovoc.europa.eu/ concepts. This dataset facilitates research in LTC, including extreme multi-label text classification, few-shot, and zero-shot learning, with documents tagged with an expansive set of descriptors.
Tuggener et al. (2020) introduce LEDGAR, a multi-label corpus of legal provisions from contracts scraped from the U.S. Securities and Exchange Commission’s website. The dataset includes over 846,000 provisions across 60,540 contracts, with an extensive label set suitable for text classification and legal studies, aiding in developing advanced legal NLP models.
Chalkidis et al. (2021) present MULTI-EURLEX, a multilingual dataset containing 65,000 EU laws translated into 23 official EU languages, annotated with EUROVOC labels. The dataset emphasizes temporal concept drift by adopting chronological splits, enhancing its utility for sophisticated LTC tasks requiring understanding nuanced legal terms across different time periods.
Papaloukas et al. (2021) introduce the Greek Legal Code dataset, categorizing approximately 47,000 Greek legislative documents into a detailed multi-level classification system. The dataset is structured into volumes, chapters, and subjects, each containing diverse legal documents from Greek legislation history, supporting LTC in the Greek legal domain.
Song et al. (2022) introduce POSTURE50K, a legal dataset containing 50,000 U.S. legal opinions annotated with Legal Procedural Postures ranging from common to rare motions. The dataset includes an innovative split strategy to support supervised and zero-shot learning evaluations, ensuring infrequent categories are adequately represented, enhancing model generalizability and testing accuracy.
Graham et al. (2023) develop a domain-specific dataset for LTC focusing on deontic modalities in contract sentences. They manually annotated contract sentences to train models for identifying deontic sentences like permissions, obligations, and prohibitions. The corpus, derived from the Contract Understanding Atticus Dataset (CUAD) (Hendrycks et al., 2021), provides a resource for studying functional categories crucial for legal analysis.
4.4. Legal Document Sumarrization
4.4.1. Task
LDS is a specialized branch of automatic summarization that focuses on condensing legal texts, such as court judgments, into clear and informative summaries. Unlike general text summarization, which extracts key details without following specific formatting rules, LDS must account for the distinct structure and specialized content of legal documents. These documents often include complex details like article numbers, statutory language, and citations that are critical for presenting the legal arguments and decisions accurately. The natural complexity of legal texts, characterized by their extensive length and detailed internal structures such as sections, articles, and paragraphs in statutes—demands tailored summarization techniques. This need is further emphasized by the hierarchical importance of documents based on their judicial origin, where the interpretation of texts can vary significantly between higher and lower court opinions (Kanapala et al., 2019).
LDS can be approached through extractive and abstractive methods. Extractive summarization techniques in LDS focus on identifying and extracting the most critical sentences or phrases directly from the text, maintaining the original wording and meaning. In contrast, abstractive summarization involves generating new sentences that paraphrase the most important information, aiming for conciseness and coherence while ensuring that the essence of the legal text is preserved. This subsection will explore existing approaches to LDS and show the small differences that set it apart from more general summarization methods and discussing the challenges and solutions specific to the legal domain.
Several systems have been specifically designed to summarize legal documents. One of the first systems in this field was the Fast Legal EXpert CONsultant (FLEXICON), created by Gelbart and Smith in 1991 (Gelbart and Smith, 1991a). FLEXICON utilizes a keyword-based approach (Gelbart and Smith, 1991b), scanning a comprehensive database of terms to pinpoint crucial segments of text. Following this, Moens et al. (1999) introduced the SALOMON system in 1999, which employs cosine similarity to cluster similar text regions, aiming to highlight relevant topics within the documents. This method aligns with other abstraction-oriented techniques seen in the work of Erkan and Radev (2004). Another system, LetSum, devised by Farzindar and Lapalme in 2004 (Farzindar, 2004), also adopts a keyword-centric strategy but uses “cue phrases” to identify text related to specific themes such as ‘Introduction’, ‘Context’, and ‘Conclusion’. Although LetSum was fairly successful in mimicking human-generated summaries, it tended to produce summaries that were excessively lengthy.
Building on previous developments in LDS, Polsley et al. (2016) introduce Casesummarizer, a tool designed for the legal domain that pre-processes legal texts into sentences, scores them using a TF-IDF matrix from extensive legal case reports, and enhances sentence scoring by identifying entities, dates, and section headings. The tool provides a user-friendly interface with scalable summaries, lists of entities and abbreviations, and a significance heat map.
Zhong et al. (2019) propose an automatic extractive summarization system for legal cases concerning Post-traumatic Stress Disorder from the US Board of Veterans’ Appeals. It employs a train-attribute-mask pipeline using a CNN classifier to iteratively select predictive sentences from case texts.
Nguyen et al. (2021) propose an RL framework to enhance deep summarization models for the legal domain, utilizing Proximal Policy Optimization with a reward function that integrates both lexical and semantic criteria. They fine-tune an extractive summarization backbone based on BERTSUM (Liu, 2019), employing a reward model that includes lexical, sentence, and keyword-level semantics to produce better legal summaries. Schraagen et al. (2022) apply an RL approach with a Bi-LSTM and a deep learning approach based on the BART transformer model to abstractive summarization of the Dutch case verdict database Rechtspraak.nl, combining extractive and abstractive summarization to retain core facts while creating concise summaries.
Zhong and Litman (2022) focus on extractive summarization of legal case decisions, proposing an unsupervised graph-based ranking model that leverages a reweighting algorithm to utilize document structure properties. They introduce a reweighting algorithm to improve sentence selection in the HipoRank model (Dong et al., 2021). It aims to reduce redundancy and enhance the selection of argumentative sentences from underrepresented sections.
Moro et al. (2023) introduce a transfer learning approach that combines extractive and abstractive summarization techniques to address the lack of labeled legal summarization datasets, outperforming previous results on the Australian Legal Case Reports dataset and establishing a new baseline for abstractive summarization.
Jain et al. (2024) propose a sentence scoring approach, DCESumm, which combines supervised sentence-level summary relevance prediction with unsupervised clustering-based document-level score enhancement. They utilize a Legal BERT-based Multi-Layer Perceptron (MLP) model to predict the summary relevance of each sentence, refining scores through deep embedded sentence clustering to enhance the selection process by considering the global context of the document.
Liu et al. (2024) present Common Law Court Judgment Summarization (CLSum), a pioneering dataset for summarizing multi-jurisdictional common law court judgments, leveraging large language models for data augmentation, summary generation, and evaluation. They employ a two-stage summarization process with techniques like sparse attention mechanisms and efficient training methods to process lengthy legal documents within limited computational resources.
4.4.2. Datasets
LDS datasets are largely built from structured court proceedings and decisions, providing rich sources for both extractive and abstractive summarization methods. These datasets often emphasize abstractive summarization to achieve concise, readable summaries that transform the original legal language into more accessible forms (Shen et al., 2022).
Zhong et al. (2019) develop a dataset from 972,522 Board of Veterans’ Appeals decisions, focusing on single-issue cases related to Post-traumatic Stress Disorder. The dataset consists of 112 carefully sampled decisions, annotated by legal experts to capture key information such as ‘Issue’, ‘Procedural History’, ‘Service History’, ‘Outcome’, ‘Reasoning’, and ‘Evidential Support’.
Shen et al. (2022) introduce Multi-LexSum, an abstractive summarization dataset tailored for U.S. federal civil rights lawsuits, containing 40,000 source documents and 9,000 expert-written summaries of diverse lengths, providing a rich resource for testing advanced summarization models.
Liu et al. (2024) publish CLSum, a dataset designed for summarizing multi-jurisdictional common law court judgments from Australia, Hong Kong, the United Kingdom, and Canada. This dataset leverages large language models for data augmentation and incorporates legal knowledge to enhance summary generation and evaluation. This dataset addresses the challenge of sparse labeled data in legal domains. CLSum includes a comprehensive collection of judgments and summaries from prominent court websites. It employs novel techniques to enrich training sets and improve model performance in few-shot and zero-shot learning scenarios.
4.5. Legal Named Entity Recognition
4.5.1. Task
NER is a fundamental task in NLP that involves identifying specific segments of text and classifying them into predefined categories such as ‘organization’, ‘person’, and ‘location’ (Layegh et al., 2023). In the legal domain, NER extends to specialized recognition tasks that focus on extracting entities unique to legal texts, such as laws, legal norms, and procedural terms. This specialized form of NER is crucial for structuring legal documents and enhancing legal IR systems. Unlike general NER systems that handle common entity types, legal NER must navigate the complex language and structured format of legal documents, underscoring the need for systems and methodologies specifically tailored to the legal context.
Dozier et al. (2010) present pioneering work in NER within legal documents such as US case law and pleadings, employing three methodologies: lookup, contextual rules, and statistical models to detect entities like judges, attorneys, and legal terms. Their system adapts these approaches to the specialized context of legal texts, processing various types of documents and extracting legal entities. This work highlights the challenges and necessary adaptations for deploying NER in the legal domain, where the specialized language and high accuracy are required for successful legal analysis.
Păis et al. (2021) develop a NER model for the legal domain that integrates Bi-LSTM cells and a Conditional Random Fields (CRFs) layer, utilizing multiple data sources and embedding types. Their system architecture combines word embeddings from pre-trained models, character embeddings, gazetteer resources from the GeoNames [^6] https://www.geonames.org/ database and JRC-Names (Steinberger et al., 2011), and known affixes to enrich the model’s understanding of legal text. Their training process involved fine-tuning the word embeddings, while dynamically learning character embeddings with subsequent Bi-LSTM layers, enhancing the model’s capability to generalize across unseen texts. They implemented the system using a modified version of NeuroNER (Dernoncourt et al., 2017), allowing for online model serving and incorporating features like dropout for regularization and gradient clipping to handle exploding gradients. They also explored ensemble methods to improve model accuracy by combining results from multiple model configurations, measuring the effectiveness through precision, recall, and F1 scores against a gold standard corpus.
Smădu et al. (2022) explore domain adaptation in Legal NER, focusing on the Romanian and German languages. They utilize a model architecture that integrates a pre-trained BERT (Devlin et al., 2019) layer for feature extraction with Bi-LSTM networks to handle sequence dependencies and CRFs for sequence tagging. Their approach employs domain adaptation techniques through a gradient reversal layer connected to a domain discriminator, aimed at reducing domain-specific biases and enhancing feature transferability across domains. This model trains on both legal and general domains simultaneously, adapting to the peculiarities of each through adversarial learning that modifies the learning process dynamically. This setup allows for improved generalization of the NER system across varied linguistic and domain contexts, though the actual performance varied with minimal improvements noted for German and decreased performance for the Romanian legal dataset.
4.5.2. Datasets
Leitner et al. (2020) present a dataset for NER focused on German federal court decisions, containing approximately 67,000 sentences and over two million tokens. This dataset features 54,000 manually annotated entities distributed across 19 fine-grained semantic classes, specifically tailored to the legal domain, such as court, judge, lawyer, law, person, and legal literature, along with over 35,000 TimeML-based time expressions. The annotations cover both broad categories like location, person, and organization, and more specialized ones like legal norms and case-by-case regulations, distinguishing between different types of legal acts and literature. This dataset’s comprehensive annotation process involved multiple cycles to refine the tagging guidelines and enhance annotation quality.
Păis et al. (2021) introduce the LegalNERo corpus, a manually annotated resource for NER in the Romanian legal domain, featuring 370 legal documents annotated with five general entity types: person, location, organization, time expressions, and legal references. This corpus was developed to support both specific legal domain NER tasks and more general NER applications by enabling compatibility with existing general-purpose NER systems. The corpus includes rich entity annotations, with legal references showing the highest token count per entity, indicating their complexity and length. The detailed annotation process, including inter-annotator agreement assessed by Cohen’s Kappa, and the subsequent mapping of entities to RDF format, highlights the corpus’s utility and precision for advancing NER research and applications within the legal domain.
Au et al. (2022) introduce the E-NER dataset, an annotated collection derived from the US SEC’s EDGAR filings, designed for legal NER. This dataset contains filings that are rich in text, such as quarterly reports (Form 10-Q) and significant event announcements (Form 8-K), from which sentences were extracted and annotated with seven named entity classes more tailored to legal content than those in the standard CoNLL dataset(Tjong Kim Sang and De Meulder, 2003). The entities include Person, Location, Organization, Government, Court, Business, and Legislation/Act, adjusting the CoNLL classes to better suit legal documents. E-NER contains significantly longer sentences compared to CoNLL and includes detailed annotations of financial entities from legal company filings.
Kalamkar et al. (2022) present a comprehensive corpus aimed at enhancing legal NER, containing 46,545 entities across 14 types identified in Indian High Court and Supreme Court judgments. This corpus, split into preamble and judgment sections, includes diverse entity types detailed in their legal NER taxonomy, such as court, petitioner, respondent, and statute, among others. The training set, drawn from judgments between 1950 and 2017, features 29,964 entities and the development and test sets, spanning 2018 to 2022. This dataset not only facilitates training and evaluation of NER models specific to the legal domain but also provides a structured framework for assessing the performance of NER systems on legal texts. Their approach leverages a combination of manual annotation and ML techniques to ensure the precision of entity recognition in legal judgments.
4.6. Large Legal Corpora
The foundational step in training an LLMs is the use of extensive corpora. For the development of a sophisticated LLMs that effectively addresses a wide range of legal NLP tasks, it is crucial to have access to large-scale legal corpora. These corpora must meet several critical criteria to ensure their effectiveness and ethical utility. First, they should be transparent in their sourcing and composition, allowing users to understand the origins and types of included data. Additionally, the privacy of individuals should be safeguarded, preventing any potential invasion of personal data. It is also imperative to minimize toxicity and bias within the corpora to promote fairness and accuracy in model outcomes. By following these principles, we can enhance the capabilities of LLMs in the legal domain. This ensures they are both powerful and reliable tools for legal analysis. In this section, we will explore the existing large legal corpora and online databases.
Zheng et al. (2021) introduce the CaseHOLD dataset, a novel benchmark for evaluating NLP models in the legal domain, designed to address the challenge of identifying the legal holdings from case texts. The dataset contains over 53,000 multiple choice questions derived from U.S. case law citations, where each question requires the identification of the correct holding from a set of potential answers. This task, simulating a fundamental lawyering skill taught in law school, involves contextual understanding and application of legal rules to factual situations. CaseHOLD is aimed at enhancing model training by focusing on semantic matching and the ability to discern nuanced legal principles. The dataset is structured to provide a challenging yet accessible resource for NLP researchers, with a clear focus on promoting deeper understanding and application of legal rules in automated systems. It utilizes a format where a cited text serves as a prompt with five answer options—one correct holding and four closely related incorrect holdings—to refine the models’ abilities to accurately reflect legal reasoning.
Chalkidis et al. (2022a) introduce the Legal General Language Understanding Evaluation (LexGLUE) benchmark, a comprehensive suite of datasets aimed at assessing the capabilities of NLP models across various legal tasks. The benchmark covers datasets such as ECtHR (Chalkidis et al., 2019a), SCOTUS [^7] https://www.supremecourt.gov, EUR-LEX, LEDGAR (Tuggener et al., 2020), UNFAIR-ToS (Lippi et al., 2019), and CaseHOLD (Zheng et al., 2021), each chosen for its complexity, relevance, and need for legal expertise. These datasets cover a range of tasks from multi-label and multi-class classification to multiple-choice questions and are split chronologically into training, development, and test sets to provide standardized evaluation metrics. For instance, ECtHR datasets focus on violations of European Convention of Human Rights provisions, SCOTUS database classifies U.S. Supreme Court opinions by legal issues, EUR-LEX database involves labeling EU laws with EuroVoc concepts, LEDGAR classifies provisions of U.S. contracts, UNFAIR-ToS identifies unfair terms in online service agreements, and CaseHOLD involves answering questions about legal rulings. This benchmark facilitates the testing of NLP models, addressing the challenges of legal text comprehension and understanding required for effective application in the legal domain.
Chalkidis et al. (2022b) present FairLex, a benchmark suite consisting of four legal datasets—ECtHR (Chalkidis et al., 2019a), SCOTUS, FSCS, and CAIL (Xiao et al., 2018)—that address the fairness of NLP applications across diverse legal jurisdictions and languages, including English, German, French, Italian, and Chinese. These datasets, curated from European Council, USA, Switzerland, and China, cover various legal tasks such as judgment prediction, issue area classification, and crime severity prediction, aiming to test the performance and fairness of LMs in recognizing and classifying legal texts. FairLex focuses on ensuring demographic, regional, and legal topic fairness by analyzing attributes like gender, age, region of origin, and legal areas within cases. Each dataset in FairLex provides a substantial number of cases, systematically divided into training, development, and test sets, and includes detailed attributes like the defendant state in ECtHR, decision direction in SCOTUS, legal areas in FSCS, and demographic details in CAIL. The CAIL dataset from China contains over a million cases focusing on criminal law, annotated with demographics and regional classifications, which are used to explore the crime severity prediction task.
Henderson et al. (2022) introduce the ‘Pile of Law’, the first and an important large corpus in the legal domain, containing a 256GB dataset of open-source English-language legal and administrative data. This dataset includes contracts, court opinions, legislative records, and administrative rules, curated to explore data sanitization norms across legal and administrative settings and serve as a tool for pre-training legal-domain LMs. They emphasize the legal norms governing privacy and toxicity filtering, detailing how the dataset reflects these norms through built-in filtering mechanisms in the collected data, which include court filings, legal analyses, and government publications. By analyzing how legal and administrative entities handle sensitive information and potentially offensive content, the paper provides actionable insights for researchers to improve content filtering practices before pre-training LLMs, thereby enhancing the ethical use of NLP in legal applications.
Rabelo et al. (2022) summarize the 8th Competition on Legal Information Extraction and Entailment (COLIEE 2021), which featured five tasks across case and statute law, engaging participants from various teams to apply diverse NLP approaches. The competition tasks included case law retrieval and entailment, as well as statute law retrieval and entailment with and without prior retrieved data. Specifically, Task 1 focused on extracting relevant supporting cases from a corpus, while Task 2 involved identifying paragraphs from cases that entail a given new case fragment. For statute law, Tasks 3 and 4 entailed retrieving and answering questions based on civil code statutes, with Task 5 challenging participants to answer without pre-retrieved statutes. The datasets used varied in complexity, from 4415 case files in Task 1 with a need to identify noticed cases without relying on citations, to the civil code-based Tasks 3, 4, and 5 which adapted to recent legal revisions in Japanese law and excluded untranslated parts, reflecting the ongoing evolution and challenge in legal NLP applications.
Barale et al. (2023) present AsyLex, a pioneering dataset tailored for Refugee Law applications, featuring 59,112 documents from Canadian refugee status determinations spanning from 1996 to 2022. This dataset is designed to enhance the capabilities of NLP models in legal research by providing 19,115 gold-standard human-annotated and 30,944 inferred labels for entity extraction and LJP. Key contributions include anonymizing decision documents, employing a robust annotation methodology, and creating datasets for specific NLP tasks like entity extraction and judgment prediction. This rich corpus, with detailed annotations across 22 categories, supports complex legal NLP tasks, thereby filling the gap in resources for the legal domain.
Niklaus et al. (2023) introduce LEXTREME, a multilingual benchmark specifically designed to evaluate LMs on legal NLP tasks, a critical step given the unique challenges of legal language. Surveying legal NLP literature from 2010 to 2022, they curate 11 datasets spanning 24 languages and cover a variety of legal domains, employing datasets that only involve human-annotated texts or those with annotations derived through clear methodological frameworks. They introduce two aggregate scores to facilitate fair comparison across models: the dataset aggregate score and the language aggregate score, revealing a performance correlation with model size on LEXTREME. The benchmark consists of three task types: Single Label Text Classification, Multi Label Text Classification, and NER, using existing splits for training, validation, and testing when available, or creating random splits otherwise. This effort marks a significant advancement in testing NLP capabilities across a diverse range of legal documents and languages.
Park and James (2023) explore the creation of a Natural Language Inference dataset within the legal domain, focusing on criminal court verdicts in Korean. Their methodology includes the innovative use of adversarial hypothesis generation to challenge annotators and enhance the robustness of the dataset, supported by visual tools for hypothesis network construction. The data collection involves extracting context from verdicts and augmenting it using Easy Data Augmentation (Wei and Zou, 2019) techniques and round-trip translation to generate a dataset for training and testing Natural Language Inference models. The study highlights issues such as annotators’ limited domain knowledge and challenges in handling long contexts but provides solutions like targeted data collection and the use of gamification to boost annotator engagement and productivity.
Goebel et al. (2024) summarize the 10th Competition on Legal Information Extraction and Entailment (COLIEE 2023), featuring four tasks across case and statute law with participation from ten different teams engaging in multiple tasks. Task 1 involves legal case retrieval, requiring participants to extract supporting cases from a corpus, and Task 2 focuses on legal case entailment, identifying paragraphs that entail aspects of a new case. Task 3 and 4, based on Japanese Civil Code statutes from the bar exam, involve retrieving relevant articles and verifying statements, respectively. The competition leverages a dataset of over 5,700 case law files and introduces new query cases and test questions sourced from recent bar exams, testing the efficacy of different teams’ approaches in handling complex legal texts and hypotheses in a controlled competitive environment.
Östling et al. (2024) introduce the Cambridge Law Corpus (CLC), a comprehensive legal dataset featuring 258,146 cases from UK courts, dating from the 16th century to the present. The corpus includes raw text and metadata across various court types, and is structured in XML format for ease of use and annotated for case outcomes in a subset of 638 cases. Additionally, the CLC is supported by a Python library for data manipulation and ML applications.
Niklaus et al. (2024) present the MultiLegalPile, the largest open-source multilingual legal corpus available, totaling 689GB and spanning 17 jurisdictions across 24 languages. This extensive dataset is designed to facilitate training of LLMs within the legal domain, featuring diverse legal text types including case law, legislation, and contracts, predominantly in English due to the integration of the ‘Pile of Law’ (Henderson et al., 2022) dataset. Through careful regex-based filtering from the mC4 corpus and manual reviews, the team ensures high precision in legal content selection. The corpus, efficiently compressed using XZ and formatted in JSONL, supports comprehensive NLP research and modeling, emphasizing its broad applicability in advancing legal AI technologies.
5. Legal Language Models And Methods For Legal Domain Adaptation
In the fast-moving field of NLP, LLMs have become a key tool for processing and understanding large amounts of unstructured text data. These models, initially trained on broad datasets like Wikipedia, have shown great skill across various language tasks. Building on this success, the legal technology community is increasingly interested in using these powerful models for Legal NLP applications. This involves adapting these general-domain models to legal texts and further training them on specialized legal documents. Such efforts aim to reduce the domain gap and customize the models to better understand the complex language used in legal documents. In this section, we will explore how these models are being adapted and applied within the legal domain to enhance Legal NLP applications.
In this section, following the methodology of this survey, we studied all peer-reviewed LMs or methods. However, due to the significant challenges present in the legal domain, there are many legal LMs that have not undergone peer review. Given the scarcity of adequate peer-reviewed resources, our research has focused on the investigation of, in order of priority, the peer-reviewed sources, then the most well-known and widely used non-peer-reviewed legal LMs. Despite their lack of formal peer review, these models have gained considerable attention and usage in the field.
5.1. Language Models
Chalkidis et al. (2020) present an in-depth analysis of applying BERT (Devlin et al., 2019), a pre-trained language model, in the legal domain, showcasing the need for domain-specific adaptation to enhance performance on legal NLP tasks. They explore three strategies: using standard BERT directly, further pre-training on legal corpora, and pre-training from scratch with legal-specific data. Their study found that both further pre-training and pre-training from scratch generally outperform the use of BERT directly. They introduce legal-bert, a specialized family of models optimized for legal text, which includes versions for varied computational capacities and demonstrates competitive performance with a lower environmental impact.
Xiao et al. (2021) introduce Lawformer, a Longformer-based (Beltagy et al., 2020) language model adapted for Chinese legal texts, designed to handle extensive document lengths common in legal data. Recognizing the limitation of standard PLMs with shorter token capacities, Lawformer employs a unique combination of sliding window, dilated sliding window, and global attention mechanisms to efficiently process long texts, making it suitable for legal AI tasks like judgment prediction and LQA. Pre-trained on a vast corpus of Chinese legal documents segmented into criminal and civil cases, Lawformer integrates complex sequential dependencies across tokens using these attention techniques, enhancing model performance for legal-specific tasks.
In the development of specialized NLP tools for Arabic legal texts, a model specifically tailored to the unique linguistic features of Arabic jurisprudence was designed, introducing AraLegal-BERT (Al-qurishi et al., 2022) midway through this innovation. This model enhances NLP applications within the legal field by adapting BERT (Devlin et al., 2019) technology to Arabic’s specific content needs, involving pre-training BERT from scratch using a broad range of legal documents, including legislative materials and contracts.
Colombo et al. (2024) introduce SaulLM-7B, a novel LLM specifically designed for legal text comprehension and generation, built on the 7 billion parameter Mistral (Jiang et al., 2023) architecture. This model is trained on an extensive English legal corpus, designed to meet the unique challenges of legal syntax and terms. SaulLM-7B uses a two-tier training approach: continued pre-training on a carefully curated 30 billion token legal dataset and an innovative instruction fine-tuning method, incorporating both generic and legal-specific instructions to enhance the model’s performance on legal tasks.
Shi et al. (2024) develop Legal-LM, a specialized language model tailored for Chinese legal consulting, enhanced with a KG to address domain-specific challenges such as data veracity and non-expert user interaction. The framework involves several steps: extensive pre-training on a rich corpus of legal texts integrated with a legal KG, keyword extraction and Direct Preference Optimization to refine responses, and the use of an external legal knowledge base for data retrieval and response validation. This multi-faceted approach ensures that Legal-LM not only comprehends complex legal language but also generates precise and user-aligned legal advice.
5.2. Methods for Improving In-Domain Adaptability of Legal Language Models
Li et al. (2022a) explore a novel adaptation of LMs for the legal domain by integrating domain-specific unsupervised data from public legal forums to optimize prefix domain adaptation, a parameter-efficient learning approach that trains only about 0.1% of the model’s parameters. They introduce a training methodology where a deep prompt is specifically tuned using a domain-adapted prefix from legal forums and then utilized in various legal tasks, demonstrating improved few-shot performance compared to full model tuning methods like legal-bert (Chalkidis et al., 2020). This approach significantly reduces computational overhead while maintaining or exceeding performance metrics across multiple legal tasks, suggesting an efficient and scalable model for legal NLP applications.
Mamakas et al. (2022) explore strategies for adapting pre-trained transformers to cope with the challenges of long legal texts within the LexGLUE benchmark, focusing on extending input capabilities and enhancing efficiency. They modify Longformer (Beltagy et al., 2020), originally extending up to 4,096 sub-words, to process up to 8,192 sub-words by reducing local attention window size and incorporating a global token at the end of each paragraph to facilitate information flow across longer texts. Additionally, they adapt legal-bert (Chalkidis et al., 2020) to employ TF-IDF representations to manage longer documents effectively, introducing variants like TF-IDF-SRT-LegalBERT, which deduplicates and sorts sub-words by TF-IDF scores, and TF-IDF-EMB-LegalBERT, which incorporates a TF-IDF embedding layer. These adaptations aim to combine the robust capabilities of transformers with the practical requirements of handling extensive legal documents, surpassing the performance of traditional linear classifiers while maintaining computational efficiency.
6. Open Research Challenges
Despite researchers’ efforts in the this interdisciplinary field and extensive advancements in AI techniques, Open Research Challenges (ORCs) still exist. In this section, we identify the ORCs and provide advice to overcome these challenges.
ORC1: Bias and Fairness
Bias and fairness are crucial concerns in the field of AI, especially at the intersection with the legal domain where decisions can deeply impact individuals’ lives. The scarcity of unbiased data in legal domains such as case law complicates the training of AI models, as these models often learn from historical decisions that may reflect existing human biases (Teichman et al., 2023; Edmond and Martire, 2019). This reliance on biased datasets can lead to unfair and biased outcomes in classification and prediction tasks. Addressing these issues is critical to ensure that AI-driven legal decisions uphold the standards of impartiality and fairness required for justice.
ORC2: Interpretability and Explainability
Interpretability and explainability are crucial across various applications in legal NLP, yet these aspects remain underexplored in many studies. The ability to trace and comprehend the decision-making process of AI systems is essential for identifying and mitigating biases. Transparent and understandable AI systems help build trust and ensure they are used responsibly, which is particularly important in legal contexts where decisions can significantly impact people’s lives. Improving these aspects of AI models is necessary to their ethical use, ensuring they meet the high standards of fairness required in legal proceedings.
ORC3: Transparent Annotation
Transparently annotated datasets are rare in the field of legal NLP. Often, studies mention the involvement of expert annotators for tasks such as classification, question answering, or prediction, but fail to provide details about the annotators’ backgrounds or the specific annotation processes used. To develop unbiased legal NLP systems, it is necessary to document the dataset creation process thoroughly. This includes providing detailed descriptions of the annotation procedures and the qualifications of the annotators involved, which is essential for ensuring the reliability and fairness of the systems trained on these datasets. Researchers need to prioritize transparency to build trust and allow for effective evaluation in the legal NLP field.
ORC4: Multilingual Capabilities
In legal NLP, enhancing multilingual capabilities remains an underdeveloped area. While efforts like MultiLegalPile (Niklaus et al., 2024) have begun to address this, there remains a gap in research for many languages, including but not limited to Persian and Arabic. These limitations significantly restricts the application of legal NLP across diverse legal systems worldwide, which is important for broader accessibility. Multilingual capabilities introduce unique challenges for legal NLP models, primarily due to the distinct linguistic structures of each language, which often require extensive fine-tuning to ensure accuracy and relevancy in legal contexts. Furthermore, each legal system possesses its own set of terms and document standards, which can vary dramatically from one language to another. Therefore, expanding research into these and other underserved languages is essential for making NLP tools universally applicable and effective.
ORC5: Ontology
The use of ontologies in the legal domain is relatively sparse, yet it holds considerable potential to enhance the robustness of AI methodologies. Ontology or KG can also enable the AI models to draw accurate inferences regarding the relationship between the terms and thereby better understand and process complicated legal texts. This approach could advance the capability of AI systems to handle complex legal reasoning and decision-making processes. However, utilizing ontology in legal NLP faces unique challenges. The complexity of legal language and the concept of ‘open texture’, where the meaning of legal terms can evolve over time, complicate the creation of static ontological models (Mommers, 2010). Legal ontologies must be dynamic, reflecting changes in law and its interpretation over time. Additionally, the integration of real-world and legal concepts within ontologies presents further complexity, as it requires accommodating both legal terms and their relevant real-world contexts (Mommers, 2010).
ORC6: Pre-processing Legal Text
Pre-processing legal texts is challenging due to the distinct nature of legal documents. Existing legal corpora are often contained of raw texts that require extensive cleaning and transformation to become suitable for ML models. Additionally, legal documents can include complex nested structures, like clauses within clauses, and cross-references to other legal cases, statutes, or provisions, making it difficult to break them into coherent units for analysis. These challenges of legal documents make it impractical to directly fine-tune LMs on these raw datasets without substantial pre-processing. This requirement complicates the use of large legal datasets, making them hard to convert into formats that NLP models can readily process and learn from. Without addressing these specific complexities, fine-tuning LMs on raw legal data becomes impractical, limiting the effectiveness of legal NLP applications.
ORC7: Reinforcement Learning from Human Feedback (RLHF)
The use of RLHF within the legal domain is notably scarce. Currently, there is only one peer-reviewed work (Nguyen et al., 2021) available that explores this approach. This indicates a significant opportunity for research and development in this area, as RLHF could potentially enhance NLP’s capability to learn and make decisions based on complex legal data under human guidance. Further exploration into this method could lead to more responsive and adaptable legal NLP systems. However, due to the complex nature of legal reasoning and the need for accurate legal knowledge in human feedback phase, RLHF’s integration into legal NLP poses some challenges. Therefore, on the human feedback side, legal experts such as lawyers and judges must provide guidance to ensure the AI accurately interprets and applies complex legal concepts.
ORC8: Expanding Legal Domain Coverage
There is a noticeable gap in the research across various areas of the legal domain, including Intellectual Property, Criminal Law, Banking Law, Family Law, and Human Rights Law. These fields have seen limited exploration across all legal NLP tasks, such as LQA and other applications. Expanding research into these areas is essential for developing comprehensive legal automated systems that can provide tailored solutions and insights highly relevant to these sectors of law.
ORC9: Small Language Models (SLMs)
Research into SLMs specific to the legal domain is notably absent. Addressing this gap could lead to more efficient, resource-conscious solutions that still maintain high performance in legal text processing and analysis. The development of SLMs tailored for legal applications could revolutionize the accessibility and scalability of legal NLP tools.
ORC10: Domain-Specific Efficient Fine-Tuning
Domain-specific efficient fine-tuning within the legal field is an underexplored area, with only two known studies (Louis et al., 2024; Li et al., 2022a) addressing it. Legal texts consist of complex structures and specialized words that standard LMs may not capture without significant adaptation. Additionally, the legal domain cover a vast array of document types, such as case law, statutes, and contracts, each requiring tailored approaches for effective model application. This diversity makes it imperative to develop fine-tuning strategies that do not only adapt a model generally but rather tailor it to understand the differences between these document types. The majority of existing approaches involve fine-tuning the entire model, which can be resource-intensive. More focused research could enable fine-tuning of legal LMs using fewer resources, enhancing the efficiency of deploying these models in practice.
ORC11: Legal Logical Reasoning
Complex legal logical reasoning remains a significant challenge in LJP, particularly in predicting prison terms. Current state-of-the-art methods struggle to achieve high accuracy in this area. This highlights a clear need for enhanced approaches that can effectively handle the complex of legal reasoning.
ORC12: Legal Named Entity Recognition
Legal NER focuses on specific challenges such as disambiguating titles, resolving nested entities, addressing co-references, managing lengthy texts, and processing machine-inaccessible PDFs. Despite its critical role in understanding and structuring legal documents, there is limited research in this area, as observed from Fig. 1.
ORC13: Stochastic Parrots
The concept of “Stochastic Parrots” pertains particularly to LLMs. It shows the concern that these models often do not truly understand language but merely mimic human patterns. This mimicry can lead to unreliable outcomes, especially in critical legal situations, if the models are not trained on high-quality, unbiased datasets. The risk is notably significant in LJP, where training on biased or unfair data could lead to irreversible outcomes, as discussed in Bender et al. (2021)’s work on the limitations of LLMs. This underscores the importance of ensuring that LLMs are trained responsibly to avoid perpetuating or amplifying existing biases in legal decisions.
ORC14: Retrieval-Augmented Generation
In the legal domain, where documents are usually lengthy, often contain cross-references, and present a variety of complex linguistic structures, LLMs can sometimes generate hallucinatory responses when faced with the task of generating accurate answers. RAG systems offer a promising solution to these challenges. RAG can mitigate issues such as the natural limitations of LLMs concerning maximum input lengths, where even extended limits may fall short due to the excessive length of many legal documents. This approach not only improves the model’s response quality but also its relevance and contextual appropriateness by incorporating more of the document’s content into the decision-making process. However, the integration of RAG into the legal domain introduces unique challenges, such as managing documents from multiple jurisdictions, ensuring temporal relevance, addressing multilingual issues, and overcoming biases in the retrieval phase. These challenges must be addressed in future research when applying RAG in the legal domain.
ORC15: Automated Legal Assistance System
To decrease human error and the costs of legal services, there’s a need for a comprehensive automated legal assistance system. This system should span all tasks within the legal NLP field, from question answering to judgment prediction, and cater to different legal specializations like civil and financial law, across multiple languages from English to Persian. Developing an accurate LLM pre-trained on a vast, diverse dataset free from biases and unfairness is crucial. This ensures that the automated legal services can reliably and equitably address a wide range of legal issues.
Summary
Table 3 illustrates the connections between ORCs and discussed areas. A direct relationship is marked with an ✓, and otherwise with a –. As shown, most ORCs are related to LJP, LQA, LTC, and LLMs, indicating more extensive research fields in these areas.
Table 3. Summary of existing ORCs in each area.
Open Research Challenges LQA LJP LTC LDS Legal NER LLMs Corpora
ORC1: Bias and Fairness
ORC2: Interpretability and Explainability
ORC3: Transparent Annotation
ORC4: Multilingual Capabilities
ORC5: Ontology
ORC6: Pre-processing Legal Text
ORC7: RLHF
ORC8: Expanding Legal Domain Coverage
ORC9: SLMs
ORC10: Domain-Specific Efficient Fine-Tuning
ORC11: Legal Logical Reasoning
ORC12: Legal NER
ORC13: Stochastic Parrots
ORC14: RAG
ORC15: Automated Legal Assistance System
7. Conclusion
Advances in AI and NLP have improved Legal NLP techniques and models. These improvements help better meet the needs of laypersons in legal matters and ease the workload of legal professionals. This survey provides a comprehensive overview of the advancements in NLP techniques used in the legal domain. Additionally, we discussed the unique characteristics of legal documents. We also reviewed existing datasets and LLMs tailored for the legal domain. Legal NER research spans multiple languages and utilizes diverse methods, from rule-based to BERT-based models. LDS has largely focused on extractive and abstractive methods, including TF-IDF and transformer-based models. In LTC, multi-class classification tasks dominate, with deep learning architectures like CNNs and Bi-LSTMs widely used. LJP primarily focuses on Chinese datasets with deep learning approaches like CNNs. LQA often leverages information retrieval techniques such as BM25, with a significant focus on statutory law. Finally, we explored key ORCs, such as the need for domain-specific fine-tuning strategies, addressing bias and fairness in legal datasets, and the importance of interpretability and explainability. Other challenges include the development of more robust pre-processing techniques, handling multilingual capabilities, and integrating ontology-based methods for more accurate legal reasoning.